Accelerator Update - July 25-July 28
Stacking and beam to Miniboone were intermittent between 8:30 AM and 11:40 AM due to a problem with a Main Injector (MI) trim power supply. The supply was eventually replaced.
Recycler (RR) studies began at 2:23 PM.
Slipstacking studies began at 2:41 PM.
Stacking, miniboone and studies were halted between 4:24 PM and 6:47 PM due to a Booster power supply tripping off. Insufficient cooling to the supply caused the trip.
Store #2817 was terminated at 7:15 PM
Switchyard studies ended at 10:50 PM.
Saturday July 26
Store #2821 was colliding by 1:55 AM with an initial luminosity of 40.44e30. We then returned to stacking.
They monitored the store and stacked for the rest of the day.
Switchyard 120 came on for studies between 4:20 PM and 7:47 PM.
Store #2821 was terminated at 11:30 PM.
Sunday July 27
At 3:07 AM Pbar returned to stacking.
The day shift started with store #2824 colliding, AntiProton Source (Pbar) stacking, and beam to Miniboone.
Lost store #2824 at 5:38 PM due to a quench at E2. The quench appeared to be caused by vacuum and/or a leaking kautzky valve. It was decided to wait until after the next store to look at these problems.
The Tev quench recovery was completed by 11:30 PM.
Monday July 28
Comments and Suggestions
|last modified 7/28/2003 email Fermilab|